# Communication-Efficient Federated Optimization over Semi-Decentralized Networks

### v2

@@ Appendix B Proof of Corollary 1
$\Psi^{K}=\frac{r_{0}}{\eta K}+a_{1}\eta+a_{2}\eta^{2}+a_{3}\eta^{2}\leq 2\left(\frac{a_{1}r_{0}}{K}\right)^{\frac{1}{2}}+2\left(\frac{\sqrt{a_{2}}r_{0}}{K}\right)^{\frac{2}{3}}+2\left(\frac{\sqrt{a_{3}}r_{0}}{K}\right)^{\frac{2}{3}}.$
@@
We mainly follow the proof of [${}^{+}$20, Lemma 17], except the first discussed case in which $\bar{\eta}$ scales on the order of $\sqrt{K}$ or $3]{K}$.
  • If $\eta=\left(\frac{r_{0}}{a_{1}K}\right)^{\frac{1}{2}}\leq\min\left\{\left(\frac{r_{0}}{a_{2}K}\right)^{\frac{1}{3}},\left(\frac{r_{0}}{a_{3}K}\right)^{\frac{1}{3}}\right\}$,
    $\Psi^{K}\leq 2\left(\frac{r_{0}a_{1}}{K}\right)^{\frac{1}{2}}+a_{2}\left(\frac{r_{0}}{a_{1}K}\right)+a_{3}\left(\frac{r_{0}}{a_{1}K}\right)\leq 2\left(\frac{r_{0}a_{1}}{K}\right)^{\frac{1}{2}}+\left(\frac{\sqrt{a_{2}}r_{0}}{K}\right)^{\frac{2}{3}}+\left(\frac{\sqrt{a_{3}}r_{0}}{K}\right)^{\frac{2}{3}},$
  • If $\eta=\left(\frac{r_{0}}{a_{2}K}\right)^{\frac{1}{3}}\leq\min\left\{\left(\frac{r_{0}}{a_{1}K}\right)^{\frac{1}{2}},\left(\frac{r_{0}}{a_{3}K}\right)^{\frac{1}{3}}\right\}$,
    $\Psi^{K}\leq 2\left(\frac{\sqrt{a_{2}}r_{0}}{K}\right)^{\frac{2}{3}}+a_{1}\left(\frac{r_{0}}{a_{2}K}\right)^{\frac{1}{3}}+a_{3}\left(\frac{r_{0}}{a_{2}K}\right)^{\frac{2}{3}}\leq 2\left(\frac{\sqrt{a_{2}}r_{0}}{K}\right)^{\frac{2}{3}}+\left(\frac{r_{0}a_{1}}{K}\right)^{\frac{1}{2}}+\left(\frac{\sqrt{a_{3}}r_{0}}{K}\right)^{\frac{2}{3}},$
  • If $\eta=\left(\frac{r_{0}}{a_{3}K}\right)^{\frac{1}{3}}\leq\min\left\{\left(\frac{r_{0}}{a_{1}K}\right)^{\frac{1}{2}},\left(\frac{r_{0}}{a_{2}K}\right)^{\frac{1}{3}}\right\}$,
    $\Psi^{K}\leq 2\left(\frac{\sqrt{a_{3}}r_{0}}{K}\right)^{\frac{2}{3}}+a_{1}\left(\frac{r_{0}}{a_{3}K}\right)^{\frac{1}{3}}+a_{2}\left(\frac{r_{0}}{a_{3}K}\right)^{\frac{2}{3}}\leq 2\left(\frac{\sqrt{a_{3}}r_{0}}{K}\right)^{\frac{2}{3}}+\left(\frac{r_{0}a_{1}}{K}\right)^{\frac{1}{2}}+\left(\frac{\sqrt{a_{2}}r_{0}}{K}\right)^{\frac{2}{3}}.$
By rearranging (5), we have
$\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[\|\nabla f(\overline{\boldsymbol{x}}^{k})\|_{2}^{2}]$
$\leq O\left(\frac{\tilde{f}}{\eta_{l}\eta_{c}T_{o}K}+\frac{\eta_{l}\eta_{c}L\sigma^{2}}{nb}+\frac{(1-p)}{(1+p)^{2}\lambda_{p}^{4}}\frac{L^{2}T_{o}^{2}\eta_{l}^{2}\eta_{c}^{2}\sigma^{2}}{b}+\frac{L^{2}T_{o}^{2}\eta_{l}^{2}\sigma^{2}}{b}+\frac{1}{nK}\mathbb{E}[\|\boldsymbol{\Phi}_{y}^{0}\|_{\mathsf{F}}^{2}]\right).$ (46)
Letting
$r_{o}=\frac{\tilde{f}}{\sqrt{1+p}\lambda_{p}},\qquad a_{1}=\frac{\sqrt{1+p}\lambda_{p}L\sigma^{2}}{nT_{o}b},\qquad a_{2}=\frac{\left(1-p\right)L^{2}\sigma^{2}}{(1+p)\lambda_{p}^{2}b},\qquad a_{3}=\frac{L^{2}\sigma^{2}}{\alpha^{2}b},$
and applying Lemma 9, then the RHS of (46) can be bounded by
$O\left(\left(\frac{L\tilde{f}\sigma^{2}}{nT_{o}bK}\right)^{\frac{1}{2}}+\left(\frac{\sqrt{1-p}L\tilde{f}\sigma}{(1+p)\lambda_{p}^{2}\sqrt{b}K}\right)^{\frac{2}{3}}+\left(\frac{L\tilde{f}\sigma}{\sqrt{(1+p)}\lambda_{p}\alpha\sqrt{b}K}\right)^{\frac{2}{3}}+\frac{1}{nK}\mathbb{E}[\|\boldsymbol{\Phi}_{y}^{0}\|_{\mathsf{F}}^{2}]\right),$
if $K$ is sufficiently large such that $\min\left\{\left(\frac{r_{0}}{a_{1}K}\right)^{\frac{1}{2}},\left(\frac{r_{0}}{a_{2}K}\right)^{\frac{1}{3}},\left(\frac{r_{0}}{a_{3}K}\right)^{\frac{1}{3}}\right\}\leq\frac{\sqrt{1+p}\lambda_{p}}{360L}$, i.e.,
$K\geq\max\left\{\frac{360^{2}nT_{o}bL\tilde{f}}{(1+p)^{2}\lambda_{p}^{4}\sigma^{2}},\,\frac{360^{3}bL\tilde{f}}{(1-p^{2})\lambda_{p}^{2}\sigma^{2}}\mathbf{1}_{p<1},\,\frac{360^{3}\alpha^{2}bL\tilde{f}}{(1+p)^{2}\lambda_{p}^{4}\sigma^{2}}\right\},$ (47)
where $\mathbf{1}_{p<1}=1$ if $p<1$ and is set to be 0 otherwise.

### v4

@@ Appendix C Proof of Corollary 1
$\Psi^{K}$ $=\frac{r_{0}}{\eta^{\prime}K}+a_{1}\eta^{\prime}+a_{2}(\eta^{\prime})^{2}\leq 2\left(\frac{a_{1}r_{0}}{K}\right)^{\frac{1}{2}}+2\left(\frac{\sqrt{a_{2}}r_{0}}{K}\right)^{\frac{2}{3}}.$
@@
We mainly follow the proof of [KLB+20, Lemma 17].
  • If $\eta^{\prime}=\left(\frac{r_{0}}{a_{1}K}\right)^{\frac{1}{2}}\leq\left(\frac{r_{0}}{a_{2}K}\right)^{\frac{1}{3}}$,
    $\Psi^{K}$ $\leq 2\bigg{(}\frac{r_{0}a_{1}}{K}\bigg{)}^{\frac{1}{2}}+a_{2}\bigg{(}\frac{r_{0}}{a_{1}K}\bigg{)}\leq 2\bigg{(}\frac{r_{0}a_{1}}{K}\bigg{)}^{\frac{1}{2}}+\bigg{(}\frac{\sqrt{a_{2}}r_{0}}{K}\bigg{)}^{\frac{2}{3}}.$
  • If $\eta^{\prime}=\left(\frac{r_{0}}{a_{2}K}\right)^{\frac{1}{3}}\leq\left(\frac{r_{0}}{a_{1}K}\right)^{\frac{1}{2}}$,
    $\Psi^{K}$ $\leq 2\left(\frac{\sqrt{a_{2}}r_{0}}{K}\right)^{\frac{2}{3}}+a_{1}\left(\frac{r_{0}}{a_{2}K}\right)^{\frac{1}{3}}\leq 2\left(\frac{\sqrt{a_{2}}r_{0}}{K}\right)^{\frac{2}{3}}+\bigg{(}\frac{r_{0}a_{1}}{K}\bigg{)}^{\frac{1}{2}}.$
    From $K\geq\max\left\{\frac{r_{0}}{a_{1}\bar{\eta}^{2}},\frac{r_{0}}{a_{2}\bar{\eta}^{3}}\right\}$, we have $\eta^{\prime}\leq\bar{\eta}$.
Then, applying Lemma 9 with $\eta^{\prime}=\alpha T_{o}\eta_{l}$ and
$r_{0}=\frac{\tilde{f}}{\sqrt{1+p}\lambda_{p}},\qquad a_{1}=\frac{\sqrt{1+p}\lambda_{p}L\sigma^{2}}{nT_{o}b},\qquad a_{2}=\frac{L^{2}\sigma^{2}}{(1+p)\lambda_{p}^{2}b},$
we bound the right hand side of (38) as
$O\left(\bigg{(}\frac{L\tilde{f}\sigma^{2}}{nT_{o}bK}\bigg{)}^{\frac{1}{2}}+\bigg{(}\frac{L\tilde{f}\sigma}{(1+p)\lambda_{p}^{2}\sqrt{b}K}\bigg{)}^{\frac{2}{3}}+\frac{1}{nK}\mathbb{E}[\|\boldsymbol{\Phi}_{y}^{0}\|_{\mathsf{F}}^{2}]\right),$
if the number of communication rounds $K$ is sufficiently large, i.e., $K\geq\max\left\{\frac{360^{2}nT_{o}bL\tilde{f}}{(1+p)^{2}\lambda_{p}^{4}\sigma^{2}},\frac{360^{3}bL\tilde{f}}{(1+p)\lambda_{p}^{2}\sigma^{2}}\right\}$ such that $\min\Big{\{}\left(\frac{r_{0}}{a_{1}K}\right)^{\frac{1}{2}},\left(\frac{r_{0}}{a_{2}K}\right)^{\frac{1}{3}}\Big{\}}\leq\bar{\eta}=\frac{\sqrt{1+p}\lambda_{p}}{360L}$.